solve each equation.
0.1(t + 0.5) + 0.2t = -0.3(t - 0.4)

Answer by TimothyLamb(4379)

0.1(t + 0.5) + 0.2t = -0.3(t - 0.4)
0.1t + 0.05 + 0.2t = -0.3t + 0.12
0.6t = 0.07
t = 0.7/0.6
t = 1.1666666666
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php